OpenStudy (anonymous):

how do u solve (1x/3) - 1 < (3/4)

OpenStudy (anonymous):

3*11=33 4y*14=56y 4y*11=44y for denominator now we have [(33+56y)/44y]>(19/16) time and cross 17*(33+56y)>44y*19 528+896y > 896y-836y>-528 60y >-528 y >-528 /60 y >-44/5

OpenStudy (anonymous):

x/3<3/4+1 x<7/4*3 x<21/4